1.Consider the following rxn: ("=" sign for rxns implies equilibrium)
2H2S = 2H2 + S2 ; Kc= 1.6x10 -2
Calculate Kc for 1/2 S2 + H2 = H2S
1
Expert's answer
2017-10-29T11:53:07-0400
For the first reaction: Kc1 = (〖〖[H〗_2]〗^2∙〖[S〗_2])/〖[H_2 S]〗^2 = 1.6x10-2 = 0.016
For the second: Kc2 = ([H_2 S])/(〖[H〗_2]∙〖〖[S〗_2]〗^(1/2) ) = √(1/K_c1 ) = √(1/0.016) = 7.9
Answer:
Kc = 7.9